# C: A Coverage Path Planning Algorithm for Unknown Environments using Rapidly Covering Graphs

### v1

@@ I-B1 Escaping Dead-ends
C provides an effective strategy for escaping from dead-ends. It is possible that the robot reaches a point where there is no unvisited node on the RCG in the local neighborhood, known as a dead-end. To escape from the dead-ends, C utilizes a dead-end escape strategy, which is built upon a concept of retreat nodes. The retreat nodes are all unvisited nodes on the RCG which are adjacent to the visited nodes. The dead-end escape strategy searches for the nearest retreat node using the A* algorithm [39] and sets it as the next waypoint. Upon reaching this node, the robot resumes the back-and-forth coverage of the remaining area.
@@ I-B2 Prevention of Coverage Holes
C provides a proactive strategy for preventing the formation of coverage holes.
Definition I.1 (Coverage Hole).
An obstacle-free uncovered region is said to be a coverage hole if it is surrounded by a) obstacles and b) covered regions.
By following a certain coverage pattern (e.g., back-and-forth or spiral), it is possible that the robot sometimes bypasses uncovered secluded regions that are surrounded by obstacles and covered regions. If left unattended, these regions form coverage holes, as shown in Fig. 3(a). Most CPP methods cover such holes later by return trajectories from distant regions, thus causing unnecessary overlaps, longer trajectory lengths and higher coverage times. To solve this issue, C detects such secluded regions in situ and covers them using TSP-based locally optimal coverage trajectories before resuming the back and forth search to cover the remaining area, as shown in Fig. 3(b). This proactive strategy prevents the formation of coverage holes, thereby improves the coverage performance and provides user satisfaction and convenience.
@@ I-C Comparison of C with the Baseline Algorithms
Table I presents the salient features of C in comparison to seven different baseline algorithms. First, the baseline algorithms are grid-based, but C is sample-based. While the baseline algorithms maintain a fine grid map, C, on the other hand, uses the RCG to represent the environment and monitor the coverage progress. The RCG is incrementally built using progressive sampling of the observed area. Second, the waypoint generation of baseline algorithms is myopic, i.e., within the local neighborhood on a grid. On the other hand, the RCG is built with a sparse-graph structure using efficient sampling and pruning techniques; thus, it provides high computational and memory efficiency with non-myopic waypoint generation. Third, the baseline algorithms could produce coverage holes, which are covered by return trajectories, thus causing strongly overlapped trajectories and high coverage times. In contrast, C actively detects potential coverage holes during navigation, performs TSP-based optimal coverage of such regions, and then resumes the back-and-forth motion. Finally, the baseline algorithms that yield spiral paths are limited in their application when turning is regarded expensive; in contrast, C produces the desired back-and-forth coverage pattern. Furthermore, in comparison to the cellular decomposition methods [2], C does not require to detect the critical points on the obstacles. Moreover, the algorithmic simplicity and computational efficiency of C make it easy to implement on robotic processors for real-time applications.
@@ I-D Contributions
The main contributions of this paper are summarized below.
  • Development of a novel CPP algorithm, called C, for real-time adaptive coverage of unknown environments, with the following features:
    • Progressive sampling to create frontier samples during navigation as the environment is discovered.
    • Incremental construction of the RCG using the frontier samples, to track the coverage progress and plan the coverage trajectory using non-myopic waypoints.
    • Escape from the dead-ends using the concept of retreat nodes.
    • Prevention of coverage holes by using TSP-based local trajectory optimization.
  • Comparative evaluation of the performance of C with the existing methods using high-fidelity simulations.
  • Testing and validation of C by real experiments on autonomous robots in a laboratory setting.
  • Comparative evaluation of the performance of C on two different CPP applications using 1) energy-constrained robots and 2) multi-robot teams.
@@ I-E Organization
The remaining paper is organized as follows. Section II presents the CPP problem and the objectives of C. Section III describes the details of the C algorithm with its iterative steps. Section IV provides the algorithm complexity and complete coverage proof. Section V shows the comparative evaluation results of C with the baseline algorithms along with the experimental validation. This section also presents the application and comparative evaluation of C for energy-constrained CPP and multi-robot CPP problems. Finally, Section VI concludes the paper with suggested recommendations for future work.
@@ II Problem Description
  • A localization device (e.g., GPS, IMU, wheel encoder, and an Indoor Localization System) or SLAM [18] to locate the position of the robot,
@@
Let $\mathcal{A}\subset\mathbb{R}^{2}$ be the unknown area populated by obstacles of arbitrary shapes that is tasked for coverage by robot $\mathcal{R}$. Let $\mathcal{A}_{o}\subset\mathcal{A}$ be the space occupied by obstacles. Then the obstacle-free space $\mathcal{A}_{f}=\mathcal{A}\setminus\mathcal{A}_{o}$ is the coverage space that is desired to be covered by the coverage device of robot $\mathcal{R}$ to perform a certain task (e.g., cleaning and demining). It is assumed that the coverage space is connected, i.e., any point in the coverage space is reachable by the robot from any other point, to form the complete coverage trajectory.
@@
Let $\lambda:[0,T]\rightarrow\mathcal{A}_{f}$, where $\lambda(t)\in\mathcal{A}_{f}$ is the point visited by robot $\mathcal{R}$ at time $t\in[0,T]$ and $T\in\mathbb{R}^{+}$ is the total coverage time. Then, the coverage trajectory is defined as $\Lambda_{T}\triangleq[\lambda(t)]_{t=0}^{T}$, with $\lambda(0)$ and $\lambda(T)$ as its start and end points, respectively.
Fig. 4(b) shows an example of the coverage trajectory and the area covered by the robot’s coverage device during navigation. The coverage trajectory has to be generated such that the robot’s coverage device covers the coverage space $\mathcal{A}_{f}$.
@@
Let $\mathcal{A}_{c}(\lambda(t))\subset\mathcal{A}_{f}$ be the area covered by the coverage device of robot $\mathcal{R}$ at point $\lambda(t)\in\Lambda_{T}$ on the coverage trajectory. Then, the coverage of $\mathcal{A}_{f}$ is said to be complete if
$\mathcal{A}_{f}\subseteq\bigcup_{t\in[0,T]}\mathcal{A}_{c}(\lambda(t)).$ (1)
The objective of C is to generate the coverage trajectory $\Lambda_{T}$ of robot $\mathcal{R}$ to provide complete coverage of $\mathcal{A}_{f}$. It is essential that C does not get stuck in dead-end situations and provide uninterrupted coverage. Furthermore, it is envisioned that C produces the desired back and forth coverage pattern with synergistic adaptations to the TSP-based locally optimal coverage trajectories to prevent the formation coverage holes. Note that significant coverage time could be saved by preventing coverage holes to avoid return trajectories from distant regions to cover such holes later. Moreover, it is desired that C minimizes the coverage time by minimizing the trajectory length, overlaps, and the number of turns.
@@ III C Algorithm
    Navigation, discovery and coverage: In this step, the robot moves towards the current waypoint of the coverage trajectory, discovers and maps the environment using its onboard sensors (e.g., the range detector) and in parallel performs the coverage task using its coverage device.
@@
    Progressive RCG growth: Once the sampling is done, this step grows the current RCG using the new samples via efficient expansion and pruning strategies.
@@ III-A Overview of the Algorithm Steps
Let the iterations of C be denoted by $i\in\{0,1...m\}$, where $m\in\mathbb{N}^{+}$. The total number of iterations is $m+1$ in which complete coverage is achieved. In each iteration $i$ the above four steps are performed. Let $p_{0}\in\mathcal{A}_{f}$ be the starting point.
@@
At iteration $i=0$, C initializes when the robot arrives at $p_{0}$ at time $t_{0}=0$, i.e., $\lambda(t_{0})=p_{0}$. Then, it stays at $p_{0}$ and senses the environment within the FOV of its range detector. Next, it samples the discovered area and generates the first installment of samples. (Details of sampling are in Section III-B2). Using these samples, it creates the first portion of RCG by setting samples as nodes and connecting nodes to form edges. (Details of RCG construction are in Section III-B3). Then, this first portion of RCG, in turn, selects a node as an intermediate goal and sets its position as the first waypoint $p_{1}\in\mathcal{A}_{f}$. (Details of waypoint generation are in Section III-B4). This finishes iteration $i=0$ with the output to the robot as waypoint $p_{1}$.
@@
Next, at iteration $i=1$, the robot moves from $p_{0}$ to $p_{1}$, thus forming the first portion of coverage trajectory. During this motion, the robot continuously discovers the environment using its range detector to incrementally build the map. While navigating the robot also performs the coverage task along the trajectory using its coverage device/sensor. The robot reaches $p_{1}$ at time $t_{1}\in(0,T]$, i.e., $\lambda(t_{1})=p_{1}$. Upon reaching $p_{1}$, it repeats Steps $2-4$ of sampling, RCG growth and trajectory generation. This finishes iteration $i=1$ with the output to the robot as the next waypoint $p_{2}\in\mathcal{A}_{f}$.
@@
In general, at iteration $i\in\{1,...m-1\}$ the robot starts at point $p_{i-1}\in\mathcal{A}_{f}$. The input to iteration $i$ is the next waypoint $p_{i}\in\mathcal{A}_{f}$, which was generated as the output of iteration $i-1$. Thus, at iteration $i$, the robot moves from $p_{i-1}$ to $p_{i}$ and forms the $i^{th}$ portion of the coverage trajectory. During this motion it discovers the environment using its range detector. It also performs the coverage task along the coverage trajectory using its coverage device/sensor. The robot reaches $p_{i}$ at time $t_{i}\in(0,T]$, i.e., $\lambda(t_{i})=p_{i}$. Upon reaching $p_{i}$, it generates samples in the newly discovered area. Note that the newly discovered area is all the area that has been sensed during motion from $p_{i-1}$ to $p_{i}$ and that has not been previously observed. Once the samples are created, it grows the existing RCG using these samples. Finally, the updated RCG generates the next waypoint $p_{i+1}\in\mathcal{A}_{f}$ to continue forming the coverage trajectory until complete coverage is achieved. Note that it is possible that at a certain iteration no new area is discovered depending on the scenario; thus, no new samples are created and RCG is not grown. In this case, the RCG relies on its existing structure and uses its available nodes to generate the next waypoint.
@@
At iteration $m-1$, the RCG generates the final destination point $p_{m}\in\mathcal{A}_{f}$. This means that at the end of iteration $m-1$, the RCG is a) fully grown, i.e., the coverage area is fully discovered, sampled, and no more eligible samples exist that could be used to grow the RCG, and b) fully exploited, i.e., the last available node is selected to generate $p_{m}$ as the final waypoint on the coverage trajectory. At the final iteration $m$, the robot simply starts at $p_{m-1}$ and moves to $p_{m}$. At this stage, no new area is discovered; thus, no new frontier samples are created and the RCG is not grown any further. Thus, upon reaching the waypoint $p_{m}$, the coverage is complete.
@@
The waypoint sequence $\mathcal{P}=\{p_{0},p_{1},...p_{m}\}$ is the sequence of intermediate goal points $p_{i}\in\mathcal{A}_{f}$ for the robot $\mathcal{R}$ to form the coverage trajectory, where $p_{0}$ and $p_{m}$ are its start and end points, respectively.
@@ III-B1 Navigation, Discovery and Coverage
Let $\mathcal{A}_{d}(\lambda(t))\subset\mathcal{A}$ be the area scanned by the range detector when the robot is at a point $\lambda(t)\in\mathcal{A}_{f}$ at time $t$. Then, the total area discovered and mapped in the $i^{th}$ iteration is the cumulation of the area scanned from time $t_{i-1}$ to $t_{i}$, given as
$\mathcal{A}_{d,i}=\bigcup_{t\in[t_{i-1},t_{i}]}\mathcal{A}_{d}(\lambda(t)).$ (2)
This area $\mathcal{A}_{d,i}$ is used for progressive sampling and growing the RCG in the next steps. The total area discovered from the beginning of the coverage operation is given as $\mathcal{A}^{\prime}_{d}=\bigcup_{i}\mathcal{A}_{d,i}$. The remaining area $\mathcal{A}_{u}=\mathcal{A}\setminus\mathcal{A}^{\prime}_{d}$ is still unknown.
Figure 6: Growth of RCG at iteration $i$ by 1) expansion into $\mathcal{F}_{i}$ using the frontier samples and 2) pruning the inessential nodes and edges.
At iteration $m$, the robot moves from $p_{m-1}$ to $p_{m}$ and senses the environment; however, since all area has been already mapped until iteration $m-1$, no new area is discovered.

### v2

@@ I-B1 Escaping Dead-ends
It is possible that the robot reaches a point where there are no unvisited nodes on the RCG within the local neighborhood, known as a dead-end. To escape from the dead-ends, C utilizes a dead-end escape strategy, which is built upon a concept of retreat nodes. The retreat nodes are all unvisited nodes on the RCG which are adjacent to the visited nodes. The dead-end escape strategy searches for the nearest retreat node using the A* algorithm [hart1968formal] and sets it as the next waypoint. Upon reaching this node, the robot resumes the back-and-forth coverage of the remaining area.
@@ I-B2 Prevention of Coverage Holes
A coverage hole is an isolated, obstacle-free, and uncovered region that is surrounded by obstacles and covered regions on all sides. While navigating a certain coverage pattern (e.g., back-and-forth or spiral), it is possible that the robot sometimes comes across such uncovered isolated regions, which if bypassed form coverage holes, as shown in Fig. 3(a). Most CPP methods cover such holes later by return trajectories from distant regions, thus causing unnecessary overlaps, longer trajectory lengths and higher coverage times. C provides a proactive strategy to prevent the formation of such coverage holes. It detects such isolated regions in situ and covers them using TSP-based locally optimal coverage trajectories before resuming the back-and-forth search to cover the remaining area, as shown in Fig. 3(b). This proactive strategy prevents the formation of coverage holes, thereby improving the coverage performance and providing user satisfaction and convenience.
@@ I-C Comparison of C with the Baseline Algorithms
Table I presents the salient features of C in comparison to seven different baseline algorithms.
  • The baseline algorithms are grid-based and maintain a fine grid map of the environment. On the other hand, C is sample-based and represents the environment using the RCG, which is incrementally built by progressive sampling of the observed area. The RCG is a minimum sufficient graph generated by efficient sampling and pruning techniques to provide high computational and memory efficiency.
  • The baseline algorithms generate myopic waypoints, i.e., within the local neighborhood on a grid. On the other hand, C generates non-myopic waypoints by means of its sparse RCG structure to reduce the total number of algorithm iterations for complete coverage.
  • The baseline algorithms could produce coverage holes during the coverage progress, which are covered later by return trajectories. This can result in strongly overlapped trajectories and high coverage times. On the other hand, C actively detects potential coverage holes during navigation, performs TSP-based optimal coverage of such regions, and then resumes the back-and-forth motion. This not only prevents the formation of undesired coverage holes but also saves coverage time.
  • The baseline algorithms using spiral paths are limited in their applications because turning is regarded as expensive. In this regard, C produces the desired back-and-forth coverage pattern.
  • The baseline algorithms using the cellular decomposition methods [acar2002sensor] require the detection of critical points on the obstacles. C does not have any such requirement.
  • Finally, the algorithmic simplicity and computational efficiency of C make it easy to implement on onboard processors for real-time applications.
@@ I-D Contributions
The paper developed a novel sample-based CPP algorithm, called C, for real-time adaptive coverage of unknown environments. C is built upon the concept of an RCG, which is incrementally constructed during robot navigation by progressive sampling along the boundary of discovered area. By virtue of efficient sampling and pruning techniques, the RCG is constructed to be a minimum-sufficient graph, which i) provides computational and memory efficiency for real-time implementation, and ii) generates non-myopic waypoints to reduce the number of iterations for complete coverage. C features a dead-end escape strategy to escape from dead-ends using the concept of retreat nodes. In addition, C features a coverage hole prevention strategy, which adapts the back-and-forth coverage trajectory to a TSP-based locally optimal trajectory to prevent the formation of coverage holes and thereby minimize the overall coverage time. The above features make C highly efficient in generating near optimal trajectories, which is successfully demonstrated by i) extensive simulations on diverse scenarios yielding significantly better performance over several baseline methods, and ii) real-experiments.
@@ I-E Organization
The remaining paper is organized as follows. Section II presents the CPP problem and the objectives of C. Section III describes the details of the C algorithm with its iterative steps. Section IV provides the algorithm complexity and complete coverage proof. Section V shows the comparative evaluation results of C with the baseline algorithms along with the experimental validation. This section also presents the application and comparative evaluation of C for energy-constrained CPP and multi-robot CPP problems. Finally, Section VI concludes the paper with suggested recommendations for future work. Appendices -A and -B show additional results and present scalability analysis for different map sizes, respectively.
@@ II Problem Description
  • A localization device (e.g., GPS, IMU, wheel encoder, and an Indoor Localization System) or SLAM [Song_sose2015] to locate the position of the robot,
@@
Let $\mathcal{A}\subset\mathbb{R}^{2}$ be the unknown area populated by obstacles of arbitrary shapes that is tasked for coverage by robot $\mathcal{R}$. Let $\mathcal{A}^{o}\subset\mathcal{A}$ be the space occupied by obstacles. Then the obstacle-free space $\mathcal{A}^{f}=\mathcal{A}\setminus\mathcal{A}^{o}$ is the coverage space that is desired to be covered by the coverage device of robot $\mathcal{R}$ to perform a certain task (e.g., cleaning and demining). It is assumed that the coverage space is connected, i.e., any point in the coverage space is reachable by the robot from any other point, to form the complete coverage trajectory.
@@
Let $\lambda:[0,T]\rightarrow\mathcal{A}^{f}$, where $\lambda(t)\in\mathcal{A}^{f}$ is the point visited by robot $\mathcal{R}$ at time $t\in[0,T]$ and $T\in\mathbb{R}^{+}$ is the total coverage time. Then, the coverage trajectory is defined as $\Lambda_{T}\triangleq[\lambda(t)]_{t=0}^{T}$, with $\lambda(0)$ and $\lambda(T)$ as its start and end points, respectively.
Fig. 4(b) shows an example of the coverage trajectory and the area covered by the robot’s coverage device during navigation. Let $\mathcal{A}^{c}(\lambda(t))\subset\mathcal{A}^{f}$ be the area covered by the robot’s coverage device at point $\lambda(t)\in\Lambda_{T}$ on the coverage trajectory. Then, let $\mathcal{A}^{c}=\bigcup_{t^{\prime}\in[0,t]}\mathcal{A}^{c}(\lambda(t^{\prime}))$ denote the total area covered until time $t$ and $\mathcal{A}^{uc}=\mathcal{A}^{f}\setminus{\mathcal{A}}^{c}$ denote the uncovered area. The coverage trajectory has to be generated such that the robot’s coverage device covers the entire free space $\mathcal{A}^{f}$.
@@
The coverage of $\mathcal{A}^{f}$ is said to be complete in time $T$ if
$\mathcal{A}^{f}\subseteq\bigcup_{t\in[0,T]}\mathcal{A}^{c}(\lambda(t)).$ (1)
$\bullet$ Objective: The objective of C is to generate the coverage trajectory $\Lambda_{T}$ of robot $\mathcal{R}$ to provide complete coverage of $\mathcal{A}^{f}$ in minimum time. It is essential that C does not get stuck at dead-ends and provide uninterrupted coverage. Furthermore, it is envisioned that C produces the desired back-and-forth coverage pattern with synergistic adaptations to the TSP-based locally optimal coverage trajectories to prevent the formation of coverage holes. Note that significant coverage time could be saved by preventing coverage holes to avoid return trajectories from distant regions to cover such holes later. Moreover, it is desired that C minimizes the coverage time by minimizing the trajectory length, overlaps, and the number of turns. Finally, it is desired that C is scalable to multi-robot CPP problems.
@@ III C Algorithm
    Navigation, discovery and coverage: In this step, the robot moves towards the current waypoint of the coverage trajectory, discovers and maps the environment using its onboard sensors (e.g., the range detector), and in parallel performs the coverage task using its coverage device.
@@
    Progressive RCG construction: Once the sampling is done, this step grows the current RCG using the new samples via efficient expansion and pruning strategies.
@@ III-A Overview of the Algorithm Steps
Let the iterations of C be denoted by $i\in\{0,1...m\}$, where $m\in\mathbb{N}^{+}$. The total number of iterations is $m+1$ in which complete coverage is achieved. In each iteration $i$ the above four steps are performed. Let $p_{0}\in\mathcal{A}^{f}$ be the starting point.
@@
$\bullet$ Iteration $i=0$: C initializes when the robot arrives at $p_{0}$ at time $t_{0}=0$, i.e., $\lambda(t_{0})=p_{0}$. Then, it stays at $p_{0}$ and senses the environment within the FOV of its range detector. Next, it samples the discovered area and generates the first installment of samples. (Details of sampling are in Section III-B2). Using these samples, it creates the first portion of RCG by setting samples as nodes and connecting nodes to form edges. (Details of RCG construction are in Section III-B3). Then, this first portion of RCG, in turn, selects a node as an intermediate goal and sets its position as the first waypoint $p_{1}\in\mathcal{A}^{f}$. (Details of waypoint generation are in Section III-B4). This finishes iteration $i=0$ with the output to the robot as waypoint $p_{1}$.
@@
$\bullet$ Iteration $i=1$: The robot moves from $p_{0}$ to $p_{1}$, thus forming the first portion of coverage trajectory. During this motion, the robot continuously discovers the environment using its range detector to incrementally build the map. While navigating the robot also performs the coverage task along the trajectory using its coverage device/sensor. The robot reaches $p_{1}$ at time $t_{1}\in(0,T]$, i.e., $\lambda(t_{1})=p_{1}$. Upon reaching $p_{1}$, it repeats Steps 242-4 of sampling, RCG growth, and trajectory generation. This finishes iteration $i=1$ with the output to the robot as the next waypoint $p_{2}\in\mathcal{A}^{f}$.
@@
$\bullet$ Iterations $i\in\{1,...m-1\}$: In general, the robot starts at point $p_{i-1}\in\mathcal{A}^{f}$. The input to iteration $i$ is the next waypoint $p_{i}\in\mathcal{A}^{f}$, which was generated as the output of iteration i1i-1. Thus, at iteration $i$, the robot moves from $p_{i-1}$ to $p_{i}$ and forms the $i^{th}$ portion of the coverage trajectory. During this motion it discovers the environment using its range detector. It also performs the coverage task along the coverage trajectory using its coverage device/sensor. The robot reaches $p_{i}$ at time $t_{i}\in(0,T]$, i.e., $\lambda(t_{i})=p_{i}$. Upon reaching $p_{i}$, it generates samples in the newly discovered area. Note that the newly discovered area is all the area that has been sensed during motion from $p_{i-1}$ to $p_{i}$ and that has not been previously observed. Once the samples are created, it grows the existing RCG using these samples. Finally, the updated RCG generates the next waypoint $p_{i+1}\in\mathcal{A}^{f}$ to continue forming the coverage trajectory until complete coverage is achieved. Note that it is possible that at a certain iteration no new area is discovered depending on the scenario; thus, no new samples are created and RCG is not grown. In this case, the RCG relies on its existing structure and uses its available nodes to generate the next waypoint.
@@
$\bullet$ Iteration i=m1i=m-1: The RCG generates the final destination point $p_{m}\in\mathcal{A}^{f}$. This means that at the end of iteration m1m-1, the RCG is a) fully grown, i.e., the coverage area is fully discovered, sampled, and no more eligible samples exist that could be used to grow the RCG, and b) fully exploited, i.e., the last available node is selected to generate $p_{m}$ as the final waypoint on the coverage trajectory.
$\bullet$ Iteration $m$: Finally, the robot starts at $p_{m-1}$ and moves to $p_{m}$. At this stage, no new area is discovered; thus, no new frontier samples are created and the RCG is not grown any further. Thus, upon reaching $p_{m}$, the coverage is complete.
@@
The waypoint sequence $\mathcal{P}=\{p_{0},p_{1},...p_{m}\}$ is the sequence of intermediate goal points $p_{i}\in\mathcal{A}^{f}$ for the robot $\mathcal{R}$ to form the coverage trajectory, where $p_{0}$ and $p_{m}$ are its start and end points, respectively.
@@ III-B1 Navigation, Discovery and Coverage
Figure 6: RCG growth at iteration $i$ by 1) expansion into $\mathcal{F}_{i}$ using the frontier samples and 2) pruning the inessential nodes and edges.
Let $\mathcal{A}^{d}(\lambda(t))\subset\mathcal{A}$ be the area scanned by the range detector when the robot is at a point $\lambda(t)\in\mathcal{A}^{f}$ at time $t$. Then, the total area discovered and mapped in the $i^{th}$ iteration is the cumulation of the area scanned from time $t_{i-1}$ to $t_{i}$, given as
$\mathcal{A}^{d}_{i}=\bigcup_{t\in[t_{i-1},t_{i}]}\mathcal{A}^{d}(\lambda(t)).$ (2)
This area $\mathcal{A}^{d}_{i}$ is used for progressive sampling and growing the RCG in the next steps. The total area discovered from the beginning of the coverage operation is given as $\mathcal{A}^{d}_{tot}=\bigcup_{i}\mathcal{A}^{d}_{i}$. The remaining area $\mathcal{A}^{u}=\mathcal{A}\setminus\mathcal{A}^{d}_{tot}$ is still unknown.
At iteration $m$, the robot moves from $p_{m-1}$ to $p_{m}$ and senses the environment; however, since all area has been already mapped until iteration m1m-1, no new area is discovered.